OpenStudy (anonymous):

find the geometric mean of 5 and 8 if necessary give the answer in simplest radical form . please show work thank you

OpenStudy (akshay_budhkar):

Geometric mean of two number a and b is \[\sqrt{ab}\]

OpenStudy (akshay_budhkar):

so in this case the geometric mean is \[\sqrt{40}=\sqrt{4\times 10}=2\sqrt{10}\]

OpenStudy (mertsj):